OpenStudy (anonymous):

What is the solution to the linear equation? 4b + 6 = 2 – b + 4 A. b=-2 B. b=0 C. b=4 D. b=6

OpenStudy (decentnabeel):

4b + 6 = 2 – b + 4 add b on both sides 4b + 6 +b= 2 – b + 4+b 5b+6=2+4 subtract 6 on both sides 5b=0 divided 5on both sides b=0

OpenStudy (anonymous):

Thank you!!

OpenStudy (decentnabeel):

wellcome:)